dattbabykayla:

Which expression is equivalent to 34 ⋅ 3−9?

TheSmartOne:

\( 3^4 \cdot 3^{-9} = ? \) You need to use the following exponent rule \( a^b \cdot a^c = a^{b+c}\) \( a^{-d} = \frac{1}{a^d} \)

TheSmartOne:

Let me know what you get or if you'd like further clarification!